the perimeter of a rectangular garden is 304 feet. If the width of the garden is 68 feet,what is its length?
Answer by Timnewman(323)

If 2(L+w)=304
L+w=152----(1)
and
w=68-----(2)
put 2 in 1
L+68=152
L=152-68
L=84ft
the lenght is 84ft
HOPE THIS HELPS
---TIMNEWMAN